# CAUSAL INFERENCE ON PROCESS GRAPHS PART II CAUSAL STRUCTURE AND EFFECT IDENTIFICATION

### v1

@@ 1.  Introduction
Causal effect identification. Section 5 is devoted to the identification of causal effects in the frequency domain. If a causal effect in the frequency domain can generically be expressed through addition/ subtraction and multiplication/division of the entries in the spectrum, then we call that effect rationally identifiable. In this paper, we make this notion more precise and, in so doing, generalise the rational identifiability term that is known for linear Gaussian SCMs. We will conclude this section by showing that the recent latent factor half-trek criterion for linear Gaussian SCMs [3] can be applied to the process graph to decide whether a given frequency domain causal effect is rationally identifiable.
@@ 2.2. The spectral trek rule
A trek from $v$ to $w$ is a triple $\tau=(\mathrm{Left}(\tau),\mathrm{top}(\tau),\mathrm{Right}(\tau))$, where $\mathrm{top}(\tau)\in V$ is a process index, and $\mathrm{Left}(\tau)\in\mathbf{P}(\mathrm{top}(\tau),v)$ and $\mathrm{Right}(\tau)\in\mathbf{P}(\mathrm{top}(\tau),w)$ are directed paths. The set of all treks from $v$ to $w$ is denoted as $\mathrm{tr}(v,w)$.
@@
$\mathcal{S}_{v,w}$ $=\sum_{\tau\in\mathrm{tr}(v,w)}\mathcal{S}^{(\tau)}.$
@@
Furthermore, let us compute the entries $\mathcal{S}_{u,w}$ and $\mathcal{S}_{v,w}$ using the spectral trek rule. For $\mathcal{S}_{u,w}$, note that there is exactly one trek connecting $v$ and $w$, namely the path $\pi$. Regarding $\mathcal{S}_{v,w}$, the set of treks is $\mathrm{tr}(v,w)=\{v\leftarrow l\to w,v\to w\}$. Consequently, we get the following expressions for the corresponding entries in the spectrum:
@@ 4.  Spectral process graph identification
Suppose $X,Y\subset V$ and $(\Phi,\Omega)\in\mathbb{R}^{\mathcal{G}}_{\mathrm{stb}}$, then $[\mathcal{S}(\Phi,\Omega)]_{X,Y}$ has rank less than or equal to $r$ if and only if there exist subsets $Z_{X},Z_{Y}\subset V$ for which $|Z_{X}|+|Z_{Y}|\leq r$, and such that $(Z_{X},Z_{Y})$ t-separates $X$ from $Y$. Generically $[\mathcal{S}_{G}]_{X,Y}$ is of rank
@@ 4.1. Spectral characterisation of process-level d-separation
It is well known that for linear Gaussian SCMs conditional independence between variables are generically equivalent to $d$-separation of the corresponding nodes on the causal graph. We will here see that this observation generalises to SVAR processes. For the following, let $G=(V,D)$ be a process graph with underlying time series graph $\mathcal{G}\in\mathcal{T}(G)$. We begin by recalling the graphical notion of $d$-separation between nodes on directed graphs.
@@
For a time series graph $\mathcal{G}\in\mathcal{T}(G)$, a parameter pair $(\Phi,\Omega)\in\mathbb{R}^{\mathcal{G}}_{\mathrm{stb}}$ is said to define a stationary distribution that is faithful with respect to the process graph $G$ if for any three disjoint sets $X,Y,Z\subset V$ it holds that $X,Y$ are $d$-separated by $Z$ if and only if $\mathbb{X}_{X}$ and $\mathbb{X}_{Y}$ are conditionally independent given $\mathbb{X}_{Z}$. Otherwise $(\Phi,\Omega)$ is said to define an unfaithful distribution with respect to $G$. The above Theorem implies that the set of parameter pairs defining an unfaithful stationary distribution is a zero set with respect to the Lebesgue measure. In particular, it is generically possible to derive the Markov equivalence class [34, 47, 35] of the process graph from the spectrum of the process.
@@ 4.2. Spectral Gessel Viennot Lemma
The proof of the trek separation theorem in the original paper [49] is based on the classical Gessel-Viennot lemma [27, 22]. In the following we will see that this lemma generalises to SVAR processes represented in the frequency domain, i.e. we establish a graphical characterisation of $\det([(I-H)^{-1}]_{X,Y})\in\mathbb{R}(z)$, where $X,Y\subset V$ are equal-sized subsets. This description is essential for the proof of Theorem 3 and also for proving graphical identifiability criteria for the link functions of SVAR processes, as we will see in the next section.
@@
Suppose $X=\{x_{1},\dots,x_{r}\}$ and $Y=\{y_{1},\dots,y_{r}\}\subset V$ are equal-sized sets of process indices. A system of directed paths from $X$ to $Y$ is a collection of directed paths $\Pi=(\pi_{1},\dots,\pi_{r})$ such that for any two distinct paths $\pi_{i},\pi_{j}\in\Pi$ it holds that they do not have the same starting or end point. In particular, the system $\Pi$ gives rise to a bijection from $X$ to $Y$, which maps $x\in X$ to the endpoint $y\in Y$ of the unique path $\pi\in\Pi$ with $x$ as its starting point.
@@
A system of directed paths $\Pi$ is said to be non-intersecting if any two paths in $\Pi$ do not share a single vertex, i.e. for any distinct $\pi,\pi^{\prime}\in\Pi$ it holds that $V(\pi)\cap V(\pi^{\prime})=\emptyset$, where $V(\pi)$ is the set of vertices visited by $\pi$. In the following we denote as $\mathbf{P}(X,Y)$ the set of all systems of non-intersecting directed paths from $X$ to $Y$. The set $\mathbf{P}(X,Y)$ is finite if $G$ is acyclic but can be infinite if $G$ contains cycles. If $X=\{x\}$ and $Y=\{y\}$, then we abbreviate $\mathbf{P}(\{x\},\{y\})$ by $\mathbf{P}(x,y)$, which is simply the set of all directed paths starting at $x$ and ending at $y$. Furthermore, if $\pi$ is a path on $G$, then we denote by $\mathrm{P}(\pi)$ the set of all directed paths $\pi^{\prime}$ that visit the same vertices as $\pi$, i.e. $\pi^{\prime}\in\mathrm{P}(\pi)$ if and only if $V(\pi^{\prime})=V(\pi)$. Note that $\mathrm{P}(\pi)$ either contains the single path $\pi$ (if $G$ does not contain cycle on the vertices $V(\pi)$) or it is an infinite set. Similarly, if $\Pi=(\pi_{i})_{i\in[r]}\in\mathbf{P}(X,Y)$, then we write $\mathrm{P}(\Pi)$ for the set of all systems $\Pi^{\prime}=(\pi_{i}^{\prime})_{1\leq i\leq r}$ with $\pi_{i}^{\prime}\in\mathrm{P}(\pi_{i})$.
Lemma 5.
Let $G$ be a process graph and $X,Y\subset V$ as above, then it holds that
$\mathbf{P}(X,Y)$ $=\bigcup_{\Pi\in\mathbf{P}_{0}(X,Y)}\mathrm{P}(\Pi),$
where $\mathbf{P}_{0}(X,Y)$ is the finite set of all non-intersecting systems of directed paths $\Pi$ such that each $\pi\in\Pi$ visits every vertex at most once.
Suppose $\Pi$ is a system of directed paths, then we define $H^{(\Pi)}$ to be the product over all path functions $h^{(\pi)}$ where $\pi\in\Pi$. In order to state the following Lemma let us recall minimal cycles [37] in the context of process graphs. A cycle on $G$ is a directed path that has the same starting and end point. A cycle $\pi$ is called minimal if the starting point is visited twice but all other vertices on $\pi$ are visited exactly once. The set of minimal cycles can be equipped with an equivalence relation. To define this relation let us recall the shift operator on cycles. Suppose $\pi=x_{1}\to\cdots\to x_{n}$ is a cycle, i.e. $x_{1}=x_{n}$. Then $\mathrm{sh}(\pi)=x_{2}\to\cdots\to x_{n}\to x_{2}$ is the shifted cycle of $\pi$. So two cycles $\pi,\pi^{\prime}$ are said to be shift equivalent if there is a $k$ such that $\pi=\mathrm{sh}^{k}(\pi^{\prime})$. Finally, we define for a directed path $\pi$ the set of minimal cycles $\mathcal{C}(\pi)$: A minimal cycle $\zeta$ is an element in $\mathcal{C}(\pi)$ if and only if there is a path $\pi^{\prime}\in\mathrm{P}(\pi)$ such that $\zeta$ is sub-path of $\pi^{\prime}$, i.e. $\zeta\leq\pi^{\prime}$. Furthermore, $\mathcal{C}_{\sim}(\pi)$ shall denote the quotient set of $\mathcal{C}(\pi)$ obtained by identifying minimal cycles that are shift equivalent.
Lemma 6.
For a path $\pi$ on which every vertex is visited at most once, it holds that
$h^{\mathrm{P}(\pi)}\coloneqq\sum_{\pi^{\prime}\in\mathrm{P}(\pi)}h^{(\pi^{\prime})}=h^{(\pi)}\prod_{\zeta\in\mathcal{C}_{\sim}(\pi)}\left(1-h^{(\zeta)}\right)^{-1},$
and for a system of non-intersecting directed paths $\Pi\in\mathbf{P}_{0}(X,Y)$ it holds that
$H^{\mathrm{P}(\Pi)}\coloneqq\sum_{\Pi^{\prime}\in\mathrm{P}(\Pi)}H^{(\Pi^{\prime})}=\prod_{\pi\in\Pi}h^{\mathrm{P}(\pi)}.$
Proof.
This Lemma follows by combining Lemma 5 and using the same arguments as in the proof Proposition A.4 in [37]. ∎
@@
Let $G$ be a directed possibly cyclic process graph and $\mathcal{G}\in\mathcal{T}(G)$, and $X,Y\subset V$ two subsets of equal size. Then it holds that
$\det([I-H]^{-1}_{X,Y})$ $=\sum_{\Pi\in\mathbf{P}_{0}(X,Y)}\mathrm{sgn}(\Pi)H^{\mathrm{P}(\Pi)},$
where $\mathbf{P}_{0}(X,Y)$ is the possibly infinite set of all non-intersecting systems of directed paths from $X$ to $Y$ and $H^{(\Pi)}=\prod_{\pi\in\Pi}h^{(\pi)}$. In particular, $\det([I-H]_{X,Y}^{-1})$ is zero (considered as an element in the field of rational functions $\mathbb{R}(\Phi,z)$) if and only if every system of directed paths from $X$ to $Y$ has an intersection.
@@
Suppose $T=(\tau_{1},\dots,\tau_{n})$ is a system of treks. The right-hand side of $T$ is the system of directed paths $\mathrm{Right}(T)=(\mathrm{Right}(\tau_{i}))_{i\in[n]}$, and the left-hand side is the system $\mathrm{Left}(T)=(\mathrm{Left}(\tau_{i}))_{i\in[n]}$. Finally, the top of the system $T$ is given by the vertices $\mathrm{top}(T)=\{\mathrm{top}(v_{i})\}_{i\in[n]}$. A system of treks is said to have a sided intersection if its right or left hand side is a system of intersecting directed paths. For subsets $X,Y\subset V$ we denote by $\mathbf{T}_{0}(X,Y)$ the set of all such trek systems $T$ that satisfy $\mathrm{Right}(T)\in\mathbf{P}_{0}(\mathrm{top}(T),Y)$ and $\mathrm{Left}(T)\in\mathbf{P}_{0}(\mathrm{top}(\mathcal{T}),X)$. That means, the right and left sides are systems of non-intersecting directed paths, and each path of these two systems visits each vertex at most once.
@@
$\det([\mathcal{S}]_{X,Y})$ $=\sum_{S\subset V:|S|=r}\det([(I-H)^{-1}]_{X,S})\det([\mathcal{S}^{\mathrm{I}}]_{S})\det([(I-H^{\ast})^{-1}]_{S,Y}).$
@@ 5.  Rational identifiability of link functions
$\mathcal{A}_{y,u}$$\coloneqq\begin{cases}[(I-H)\mathcal{S}]_{y,u}&\text{if }y\in\mathrm{htr}_{L^{\prime}}(W\cup\{v\})\\ \mathcal{S}_{y,u}&\text{if }y\not\in\mathrm{htr}_{L^{\prime}}(W\cup\{v\})\end{cases},$
@@
The matrices $\mathcal{A},\mathcal{B}$ and the vector $g$ are rationally identifiable by assumption, i.e. thy can be constructed for almost all $(\Phi,\Omega)\in\mathbb{R}^{\mathcal{G}}_{\mathrm{stb}}$. To construct $f$, note that due to the assumption that $(Y,W,L^{\prime})$ satisfies the LF-HTC with respect to $v$, it holds that
$[\mathcal{S}^{\mathrm{I}}]_{Y,W\cup\{v\}}=[H]_{L^{\prime},O}^{\top}[\mathcal{S}^{\mathrm{I}}]_{L^{\prime},L^{\prime}}[H]_{L^{\prime},O}^{\ast}$
The matrix $[\mathcal{S}^{\mathrm{I}}]_{Y,W\cup\{v\}}$ hast at least rank $r$, for generic parameter configurations, but cannot have full column rank (as a matrix over $\mathbb{R}(z)$), since $|L^{\prime}|=|W|=r$ and condition 2 of the LF-HTC combined with Proposition 5 applied to the sub-graph $G^{\prime\prime}=(O\cup L,D_{L})$. So, for generic $(\Phi,\Omega)\in\mathbb{R}^{\mathcal{G}}_{\mathrm{stb}}$, there is a $f\in\mathbb{R}(z)^{W}$ such that
$[\mathcal{S}^{\mathrm{LI}}]_{Y,Z}f=[\mathcal{S}^{\mathrm{LI}}]_{Y,v}.$
@@
It remains to show that for generic parameter choices $(\Phi,\Omega)$, the matrix $\begin{bmatrix}\mathcal{A}&\mathcal{B}\end{bmatrix}$ is generically invertible. To see that, let us construct a sub-graph $G^{\prime}=(V,D^{\prime})$ of $G$. The set of links on $G^{\prime}$ are defined as follows:
$D^{\prime}$ $\coloneqq D\setminus\left\{x\to y\in D_{O}\mid y\in W\cup(Y\cap\mathrm{htr}_{H}(W\cup\{v\}\}))\right\}$
Accordingly, we remove edges from the time series graph $\mathcal{G}$ to construct a subgraph $\mathcal{G}^{\prime}=(V\times\mathbb{Z},\mathcal{D}^{\prime})$ of $\mathcal{G}$. Specifically, the set of links on $\mathcal{G}^{\prime}$ is as follows
$\mathcal{D}^{\prime}$ $\coloneqq\mathcal{D}\setminus\left\{x(t-k)\to y(t)\in\mathcal{D}_{O}\mid y\in W\cup(Y\cap\mathrm{htr}_{H}(W\cup\{v\}\}),t\in\mathbb{Z}\right\}$
Let us denote by $\mathcal{S}^{\prime}$ the parameterisation of the spectral density for the time series graph $\mathcal{G}^{\prime}$. Using the trek-rule one can show that
$[\mathcal{S}^{\prime}]_{Y,\mathrm{pa}(v)\cup W}=\begin{bmatrix}\mathcal{A}&\mathcal{B}\end{bmatrix}.$
The left side of this equation is generically invertible because there is a system of treks without sided intersection from $Y$ to $\mathrm{pa}(v)\cup W$ on the subgraph $G^{\prime}$. The existence of such a trek system follows from the fact that every latent factor half trek from $Y$ to $\mathrm{pa}(v)\cup W$ on $G$ is also a latent factor half trek on $G^{\prime}$ and the assumption that $(Y,W,L^{\prime})$ satisfies the LF-HTC (in particular condition 3). Hence, we conclude by Proposition 5 that $\mathcal{S}^{\prime}_{Y,\mathrm{pa}(v)\cup W}$ is generically invertible. This concludes the proof. ∎

### v2

@@ 1.  Introduction
Causal effect identification. Section 5 is devoted to the identification of causal effects in the frequency domain. If a causal effect in the frequency domain can generically be expressed through addition/ subtraction and multiplication/division of the entries in the spectrum, then we call that effect rationally identifiable. In this paper, we make this notion more precise and, in so doing, generalise the rational identifiability term that is known for linear Gaussian SCMs. We will conclude this section by showing that the recent latent factor half-trek criterion for linear Gaussian SCMs [3] can be applied to the process graph to decide whether a given frequency domain causal effect is rationally identifiable. However, we need to require that the process graph is acyclic, while [3] allows also for cyclic graphs.
@@ 2.2. The spectral trek rule
A trek from $v$ to $w$ is a triple $\tau=(\mathrm{Left}(\tau),\mathrm{top}(\tau),\mathrm{Right}(\tau))$, where $\mathrm{top}(\tau)\in V$ is a process index, and $\mathrm{Left}(\tau)\in\mathbf{P}(\mathrm{top}(\tau),v)$ and $\mathrm{Right}(\tau)\in\mathbf{P}(\mathrm{top}(\tau),w)$ are directed paths. The set of all treks from $v$ to $w$ is denoted as $\mathbf{T}(v,w)$.
@@
$\mathcal{S}_{v,w}$ $=\sum_{\tau\in\mathbf{T}(v,w)}\mathcal{S}^{(\tau)}.$
@@
Furthermore, let us compute the entries $\mathcal{S}_{u,w}$ and $\mathcal{S}_{v,w}$ using the spectral trek rule. For $\mathcal{S}_{u,w}$, note that there is exactly one trek connecting $v$ and $w$, namely the path $\pi$. Regarding $\mathcal{S}_{v,w}$, the set of treks is $\mathbf{T}(v,w)=\{v\leftarrow l\to w,v\to w\}$. Consequently, we get the following expressions for the corresponding entries in the spectrum:
@@ 4.  Spectral process graph identification
Let $G$ is an acyclic process graph, and $X,Y\subset V$ not necessarily disjoint subsets, and $(\Phi,\Omega)\in\mathbb{R}^{\mathcal{G}}_{\mathrm{stb}}$ an SVAR-parameter pair, then $[\mathcal{S}(\Phi,\Omega)]_{X,Y}$ has rank less than or equal to $r$ if and only if there exist subsets $Z_{X},Z_{Y}\subset V$ for which $|Z_{X}|+|Z_{Y}|\leq r$, and such that $(Z_{X},Z_{Y})$ t-separates $X$ from $Y$. Generically $[\mathcal{S}_{G}]_{X,Y}$ is of rank
@@ 4.1. Spectral characterisation of process-level d-separation
It is well known that for linear Gaussian SCMs conditional independence between variables are generically equivalent to $d$-separation of the corresponding nodes on the causal graph. We will here see that this observation generalises to SVAR processes. For the following, let $G=(V,D)$ be an acyclic process graph with underlying time series graph $\mathcal{G}\in\mathcal{T}(G)$. We begin by recalling the graphical notion of $d$-separation between nodes on directed graphs.
@@
For a time series graph $\mathcal{G}\in\mathcal{T}(G)$, a parameter pair $(\Phi,\Omega)\in\mathbb{R}^{\mathcal{G}}_{\mathrm{stb}}$ is said to define a stationary distribution that is faithful with respect to the process graph $G$ if for any three disjoint sets $X,Y,Z\subset V$ it holds that $X,Y$ are $d$-separated by $Z$ if and only if $\mathbb{X}_{X}$ and $\mathbb{X}_{Y}$ are conditionally independent given $\mathbb{X}_{Z}$. Otherwise $(\Phi,\Omega)$ is said to define an unfaithful distribution with respect to $G$. Theorem 4 implies that the set of parameter pairs defining an unfaithful stationary distribution is a zero set with respect to the Lebesgue measure. In particular, it is generically possible to derive the Markov equivalence class [34, 47, 35] of the process graph from the spectrum of the process.
@@ 4.2. Spectral Gessel Viennot Lemma
The proof of the trek separation theorem in the original paper [49] is based on the classical Gessel-Viennot lemma [27, 22]. In the following we will see that this lemma generalises to SVAR processes represented in the frequency domain, i.e. we establish a graphical characterisation of $\det([(I-H)^{-1}]_{X,Y})\in\mathbb{R}(z)$, where $X,Y\subset V$ are equal-sized subsets. This description is essential for the proof of Theorem 3 and also for proving graphical identifiability criteria for the link functions of SVAR processes, as we will see in the next section. For the rest of this subsection we assume that the process graph $G$ is acyclic.
@@
Suppose $X=\{x_{1},\dots,x_{r}\}$ and $Y=\{y_{1},\dots,y_{r}\}\subset V$ are equal-sized sets of process indices. A system of directed paths from $X$ to $Y$ is a collection of directed paths $\Pi=(\pi_{1},\dots,\pi_{r})$ such that for any two distinct paths $\pi_{i},\pi_{j}\in\Pi$ it holds that they do not have the same starting or end point. In particular, the system $\Pi$ gives rise to a bijection from $X$ to $Y$, which maps $x\in X$ to the endpoint $y\in Y$ of the unique path $\pi\in\Pi$ with $x$ as its starting point. Given the ordering of $X$ and $Y$, $\Pi$ induces a permutation $\alpha$ on the set $\{1,\dots,r\}$ where $\alpha(i)=j$ if the unique path $\pi\in\Pi$ starting at $x_{i}$ ends at $y_{j}$. In the following, we write $\mathrm{sgn}(\Pi)$ for the sign of the associated permutation $\alpha$.
@@
A system of directed paths $\Pi$ is said to be non-intersecting if any two paths in $\Pi$ do not share a single vertex, i.e. for any distinct $\pi,\pi^{\prime}\in\Pi$ it holds that $V(\pi)\cap V(\pi^{\prime})=\emptyset$, where $V(\pi)$ is the set of vertices visited by $\pi$. In the following we denote as $\mathbf{P}(X,Y)$ the set of all systems of non-intersecting directed paths from $X$ to $Y$, and for $\Pi\in\mathbf{P}(X,Y)$ we define
$H^{(\Pi)}$ $\coloneqq\prod_{\pi\in\Pi}h^{(\pi)}.$
The set $\mathbf{P}(X,Y)$ is finite if $G$ is acyclic but can be infinite if $G$ contains cycles. If $X=\{x\}$ and $Y=\{y\}$, then we abbreviate $\mathbf{P}(\{x\},\{y\})$ by $\mathbf{P}(x,y)$, which is simply the set of all directed paths starting at $x$ and ending at $y$.
@@
Let $G$ be a directed acyclic process graph and $\mathcal{G}\in\mathcal{T}(G)$, and $X,Y\subset V$ two subsets of equal size. Then it holds that
$\det([I-H]^{-1}_{X,Y})$ $=\sum_{\Pi\in\mathbf{P}(X,Y)}\mathrm{sgn}(\Pi)H^{(\Pi)}.$
In particular, $\det([I-H]_{X,Y}^{-1})$ is zero (considered as an element in the field of rational functions $\mathbb{R}(\Phi,z)$) if and only if every system of directed paths from $X$ to $Y$ has an intersection.
@@
Suppose $T=(\tau_{1},\dots,\tau_{n})$ is a system of treks. The right-hand side of $T$ is the system of directed paths $\mathrm{Right}(T)=(\mathrm{Right}(\tau_{i}))_{i\in[n]}$, and the left-hand side is the system $\mathrm{Left}(T)=(\mathrm{Left}(\tau_{i}))_{i\in[n]}$. Finally, the top of the system $T$ is given by the vertices $\mathrm{top}(T)=\{\mathrm{top}(v_{i})\}_{i\in[n]}$. A system of treks is said to have a sided intersection if its right or left hand side is a system of intersecting directed paths. For subsets $X,Y\subset V$ we denote by $\mathbf{T}(X,Y)$ the set of all such trek systems $T$ that satisfy $\mathrm{Right}(T)\in\mathbf{P}(\mathrm{top}(T),Y)$ and $\mathrm{Left}(T)\in\mathbf{P}(\mathrm{top}(\mathcal{T}),X)$. That means, the right and left sides are systems of non-intersecting directed paths, i.e. $\mathbf{T}(X,Y)$ is the set of trek systems without sided intersection.
@@
$\det([\mathcal{S}]_{X,Y})$ $=\sum_{S\subset V:|S|=r}\det([(I-H)^{-1}]_{X,S})\det([\mathcal{S}^{\mathrm{I}}]_{S})\det([(I-H^{\ast})^{-1}]_{S,Y}).$ (12)
@@
For subsets $X,Y\subset V$ of equal size $r$, the combination of equation (12) and Proposition 4 yields
$\det([\mathcal{S}]_{X,Y})$ $=\sum_{T\in\mathbf{T}(X,Y)}\mathrm{sgn}(T)\mathcal{S}^{(T)},$ (13)
where for $T=(\tau_{1},\dots,\tau_{r})\in\mathbf{T}(X,Y)$ a system of treks without sided intersection and $\mathcal{S}^{(T)}=\prod_{i=1}^{r}\mathcal{S}^{(\tau_{i})}$ the associated product of trek functions and $\mathrm{sgn}(T)$ the sign of the permutation associated with $T$.
@@ 5.  Rational identifiability of link functions
Let $T\in\mathbf{T}(X,Y)$ be a system of treks from $X$ to $Y$. Then let $\mathcal{G}(T)=(V\times\mathbb{Z},\mathcal{D}(T))$ be the time series subgraph of $\mathcal{G}$, where $v\to_{k}w$ is a causal relation on $\mathcal{G}(T)$ if and only if $v\to_{k}w$ on $\mathcal{G}$ and $v\to w$ is a link that appears in the system of treks $T$. Accordingly, $G(T)$ denotes the process graph of the time series graph $\mathcal{G}(T)$. Since $\mathcal{D}(T)$ is a subset of $\mathcal{D}$, there is a canonical embedding $\iota_{T}:\mathbb{R}^{\mathcal{D}(T)}_{\mathrm{stb}}\to\mathbb{R}^{\mathcal{D}}_{\mathrm{stb}}$. If $T\in\mathbf{T}(X,Y)$ is such that $G(T)$ acyclic, then for $\Phi\in\mathbb{R}^{\mathcal{D}(T)}_{\mathrm{stb}}$ one can apply equation (13) to get that
$\det([\mathcal{S}(\iota_{T}(\Phi),\Omega,z)]_{X,Y})$ $=\sum_{T^{\prime}\in\mathbf{T}_{G(T)}(X,Y)}\mathrm{sgn}(T)\mathcal{S}^{(T)}(\iota_{T}(\Phi),\Omega),$
where $\mathbf{T}_{G(T)}(X,Y)$ is the finite non-empty set of non-intersecting trek systems from $X$ to $Y$ on the subgraph $G(T)$. So it follows from Proposition 5 that $\det([\mathcal{S}]_{X,Y})\in\mathbb{R}(\Phi,\Omega,z)$ is non-zero.
@@
$\mathcal{A}_{y,u}$$\coloneqq\begin{cases}[(I-H)^{\top}\mathcal{S}]_{y,u}&\text{if }y\in\mathrm{htr}_{L^{\prime}}(W\cup\{v\})\\ \mathcal{S}_{y,u}&\text{if }y\not\in\mathrm{htr}_{L^{\prime}}(W\cup\{v\})\end{cases},$
@@
The matrices $\mathcal{A},\mathcal{B}$ and the vector $g$ are rationally identifiable by assumption, i.e. they can be constructed for almost all $(\Phi,\Omega)\in\mathbb{R}^{\mathcal{G}}_{\mathrm{stb}}$. To construct $f$, note that due to the assumption that $(Y,W,L^{\prime})$ satisfies the LF-HTC with respect to $v$, it holds that
$[\mathcal{S}^{\mathrm{LI}}]_{Y,W\cup\{v\}}=[H]_{L^{\prime},O}^{\top}[\mathcal{S}^{\mathrm{I}}]_{L^{\prime},L^{\prime}}[H]_{L^{\prime},O}^{\ast}$
The matrix $[\mathcal{S}^{\mathrm{LI}}]_{Y,W\cup\{v\}}$ has at least rank $r$, for generic parameter configurations, but cannot have full column rank (as a matrix over $\mathbb{R}(z)$), since $|L^{\prime}|=|W|=r$ and condition 2 of the LF-HTC combined with Proposition 5 applied to the sub-graph $G^{\prime\prime}=(O\cup L,D_{L})$. So, for generic $(\Phi,\Omega)\in\mathbb{R}^{\mathcal{G}}_{\mathrm{stb}}$, there is a $f\in\mathbb{R}(z)^{W}$ such that
$[\mathcal{S}^{\mathrm{LI}}]_{Y,W}f=[\mathcal{S}^{\mathrm{LI}}]_{Y,v}.$
@@
It remains to show that for generic parameter choices $(\Phi,\Omega)$, the matrix $\begin{bmatrix}\mathcal{A}&\mathcal{B}\end{bmatrix}$ is invertible. Let us therefore write $F\coloneqq\det(\begin{bmatrix}\mathcal{A}&\mathcal{B}\end{bmatrix})$ which we view as an element in $\mathbb{R}(\Phi,\Omega,z)$. We need to show that $F$ is non-zero. Let $\mathbf{T}_{F}$ be the set of trek systems $T^{\prime}$ such that the following conditions are satisfied:
  1. 1.
    If $\tau\in T^{\prime}$ is a trek from $y\in Y\cap\mathrm{htr}_{L^{\prime}}(W\cup\{v\})$ to $u\in\mathrm{pa}_{O}(v)$, then $\mathrm{Left}(\tau)$ does not contain any link $x\to y\in D_{O}$.
  2. 2.
    If $\tau\in T^{\prime}$ is a trek from $y\in Y$ to $w\in W$, then $\mathrm{Right}(\tau)$ does not contain any link $v\to w\in D_{O}$. If in addition $y\in Y\cap\mathrm{htr}_{L^{\prime}}(W\cup\{v\})$, then it is also required that $\mathrm{Left}(\tau)$ does not contain any link $x\to y\in D_{O}$.
If $T\in\mathbf{T}(X,Y)$ such that $G(T)$ is acyclic, then for $(\Phi,\Omega)\in\mathbb{R}^{\mathcal{G}(T)}_{\mathrm{stb}}$ it holds that
$F(\iota_{T}(\Phi),\Omega,z)$ $=\sum_{T\in\mathbf{T}_{F}\cap\mathbf{T}_{G(T)}(X,Y)}\mathrm{sgn}(T)\mathcal{S}^{(T)}(\iota_{T}(\Phi),\Omega,z),$ (15)
To see that $F$ is not the zero element in $\mathbb{R}(\Phi,\Omega)$ it suffices to identify some $(A,\Omega,z)$ such that $F$ evaluated at that triple is non-zero.
In order to find such a triple, we consider a system of latent factor half-treks $T\in\mathbf{T}(Y,W\cup\mathrm{pa}_{O}(v))$ of the form specified in Definition 8 and that additionally satisfies the properties spelled out in Lemma 6. Note that the system of latent factor half-treks $T$ is an element in $\mathbf{T}_{F}$, which follows from Definition 8. Using Lemma 6 we conclude that $\mathbf{T}_{F}\cap\mathbf{T}_{G(T)}(X,Y)=\{T\}$. For $(A,\Omega)\in\mathbb{R}^{\mathcal{G}(T)}_{\mathrm{stb}}$ equation (15) reduces to
$F(\iota_{T}(A),\Omega,z)=\mathrm{sgn}(T)\mathcal{S}^{(T)}(\iota_{T}(A),\Omega,z)=\det([\mathcal{S}_{\mathcal{G}(T)}(A,\Omega,z)])_{Y,\mathrm{pa}_{O}(v)\cup W}$
and this suffices to conclude that $F$ is non-zero as an element in $\mathbb{R}(\Phi,\Omega,z)$, and therefore that the matrix $\begin{bmatrix}\mathcal{A}&\mathcal{B}\end{bmatrix}$ is generically invertible. ∎